OpenStudy (dariyana):

what is the sqrt of 96

OpenStudy (uri):



OpenStudy (uri):

or 9.7979

OpenStudy (whpalmer4):

Ah, maybe you're wondering just how to figure this out?

OpenStudy (whpalmer4):

\[\sqrt{96} = \] You could break 96 down into its prime factorization: \[96=2*48 = 2*2*24=2*2*2*12\]\[=2*2*2*2*6=2*2*2*2*2*3\]For each pair of identical factors under square root sign, you can move one of them outside of the radical: \[\sqrt{96} = \sqrt{(2*2)*(2*2)*2*3} = 2*2* \sqrt{2*3}=4\sqrt{6}\] You could also notice that \(96 = 6*16\), and \(16\) is a perfect square (\(4^2=16)\): \[\sqrt{96} = \sqrt{6*16} = \sqrt{6}*\sqrt{16} = \sqrt{6}*4 = 4\sqrt{6}\]